# LLM-Enabled In-Context Learning for Data Collection Scheduling in UAV-assisted Sensor Networks

### v1

@@ Abstract
Unmanned Aerial Vehicles (UAVs) are increasingly being used in various private and commercial applications, e.g. traffic control, package delivery, and Search and Rescue (SAR) operations. Machine Learning (ML) methods used in UAV-assisted Sensor Networks (UASNETs) and especially in Deep Reinforcement Learning (DRL) face challenges such as complex and lengthy model training, gaps between simulation and reality, and low sample efficiency, which conflict with the urgency of emergencies such as SAR operations. This paper proposes In-Context Learning (ICL)-based Data Collection Scheduling (ICLDC) scheme, as an alternative to DRL in emergencies. The UAV collects and transmits logged sensory data, to an LLM, to generate a task description in natural language, from which it obtains a data collection schedule to be executed by the UAV. The system continuously adapts by adding feedback to task descriptions and utilizing feedback for future decisions. This method is tested against jailbreaking attacks, where task description is manipulated to undermine network performance, highlighting the vulnerability of LLMs to such attacks. The proposed ICLDC outperforms the Maximum Channel Gain by reducing cumulative packet loss by approximately 56%. ICLDC presents a promising direction for intelligent scheduling and control in UAV-assisted data collection.
@@
Unmanned Aerial Vehicles, Large Language Models, In-Context Learning, Jailbreaking Attacks, Network Edge
@@ I Introduction
Unmanned Aerial Vehicles (UAVs) enjoy controlled mobility and rapid deployment. These features allow them to be used in many civilian and commercial applications, e.g., traffic control[1], package delivery[2], and crop monitoring[3]. UAVs are also employed to collect sensory data in harsh environments, such as natural disaster monitoring, and border surveillance[4]. In particular, Search and Rescue (SAR) operations are critical endeavors in disaster management and public safety, often involving high-risk situations where human lives are at stake. In recent years, the integration of UAV technology into SAR operations has revolutionized the field, offering a powerful tool to enhance the efficiency, effectiveness, and safety of rescue missions[5]. A common trend in SAR operations is the use of Deep Reinforcement Learning (DRL) to optimize UAV-assisted data collection in disaster areas[6],[7]. However, DRL optimizers often suffer from low sample efficiency, simulation-to-reality gaps, and the need for complex model training and fine-tuning which are not compatible with the urgent and emergency nature of SAR missions[8],[9].
@@
On the other hand, Large Language Models (LLMs) are transformative technologies with successful applications in education, finance, healthcare, biology, and more[10]. Their potential extends to network management and optimization, particularly through In-Context Learning (ICL), a feature that allows LLMs to learn from language-based descriptions and demonstrations. Compared to existing methods like convex optimization or conventional Machine Learning (ML), LLM-enabled ICL offers key advantages, it relies on the LLM’s inference process, eliminating the need for complex model training and fine-tuning, which are often bottlenecks in ML. In addition, tasks can be formulated using natural language, making it accessible to operators without deep technical expertise[11].
@@
These benefits make LLM-enabled ICL a promising approach for simplifying and enhancing network management and optimization tasks[11]. For instance, UAVs can use LLM-enabled ICL to optimize data collection scheduling and dynamically adapt to changing environments. However, the use of LLMs also introduces vulnerabilities to jailbreaking attacks which refers to attempts to bypass or undermine the safety filters and restrictions integrated into LLMs. The main objective is to circumvent the model built-in limitations, enabling it to perform actions or produce outputs that are typically restricted or prohibited.
Figure 1: Overall design of the proposed ICLDC.
@@
Given the challenges and opportunities of LLMs, this paper proposes ICL-based Data Collection Scheduling (ICLDC) scheme. Specifically, we conduct ICL to optimize data collection scheduling in UAV-assisted Sensor Networks (UASNETs), the UAV interacts with the environment to collect logged sensory data and transmit it to the LLM for analysis and decision-making. The LLM is initialized with a robust system prompt that defines its role, constraints, and ethical guidelines, ensuring it adheres to safe and secure operation principles. In particular, the LLM begins by crafting a detailed task description in natural language, which includes the task goal, input data, rules, examples, output, and feedback. Then, the LLM leverages this information to generate a data collection schedule, and the UAV executes it. Following this, the data collection schedule made by the LLM, along with the environment state (e.g., queue length, and channel condition of ground sensors) and the resulting system performance (e.g., packet loss), is recorded as feedback and added to the task description. After that, when the next environment state arises, feedback is used to guide the LLM in making subsequent decisions, enabling continuous adaptation and improvement. Under attack, task description is manipulated to impact the network cost negatively.
@@
  • ICLDC, an LLM integrated with ICL to optimize data collection scheduling in UASNETs. In contrast to conventional methods such as DRL, which require extensive model training and fine-tuning, the proposed ICLDC enables the UAV to make scheduling decisions based on natural language descriptions and demonstrations. This approach is instrumental in dynamic environments such as disaster monitoring or SAR missions.
@@
The rest of this paper is organized as follows: Section II discusses background concepts on data collection scheduling, ICL, and jailbreaking attacks. Section III presents the literature review. Section IV presents the system model comprising problem formulation and communication protocol. Section V presents the attack model with a detailed procedure. Section VI presents the proposed approach. Section VII devotes to numerical results and discussions and finally section VIII concludes the paper.
@@ II Background Concepts
This section introduces some background concepts related to data collection scheduling in UASNETs, ICL, and jailbreaking including attack types and template completion attacks.
II-A UAV Data Collection
UASNETs leverage the unique capabilities of UAVs to operate in SAR missions and remote environments, making them ideal for aerial data collection. UAVs can serve as Aerial Base Stations (ABS) or relays, extending the coverage and connectivity of sensor networks. In UASNETs, the UAV is employed to hover over the target area, moving sufficiently close to ground sensors, exploiting short-distance LoS communication links for data collection. However, selecting a ground sensor for data collection may lead to buffer overflows at other sensors, if those sensors buffers are already full while new data keeps arriving. Moreover, the transmissions of ground sensors that are far away from the UAV and experience poor channel conditions are prone to errors at the UAV. Adequately scheduling data collection is critical to prevent data queue overflow and communication failure[12],[13].
II-B In-Context Learning
Table I: Summary of the Attack Model
Impact Increased packet loss.
II-C Jailbreaking Attacks
A prominent black-box method, template completion attacks, disguises harmful queries within structured templates, compelling LLMs to complete dangerous responses unknowingly. Attackers achieve this through scenario nesting (framing queries within fictional or role-playing contexts), context-based manipulation (using misleading continuity to elicit unauthorized outputs), and code injection (embedding malicious instructions within programming tasks). These attacks are particularly dangerous as they evade keyword-based filters, exploit the LLM’s pattern recognition capabilities, and work effectively even against well-aligned models[15]. Table II provides a summary of jailbreaking attack in LLMs.
III Related Work
When searching the current literature for the use of LLM-enabled ICL in the design, deployment, and operation of wireless systems and UAV communications, we realized that this topic is rarely investigated, particularly in the context of UAV communications and security. This section briefly mentions the related work we found and its focus.
Hao et al. [11] propose an LLM-enabled ICL algorithm for base station transmission power control that can handle both discrete and continuous state problems, their approach outperforms state-of-the-art DRL algorithms. In another work, Hao et al. [14] provide a detailed investigation of different prompting techniques such as ICL, chain-of-thought, and self-refinement, and propose novel prompting schemes for network optimization, and the case studies confirm their satisfactory performance. Dong et al. [16] present a comprehensive survey on ICL where advanced techniques, comprising training and prompt designing strategies, are discussed. Zheng et al. [17] investigate the potential of deploying LLMs at the 6G edge, identify the critical challenges for LLM deployment at the edge, and envision the 6G Mobile Edge Computing (MEC) architecture for LLMs. Liu et al. [18] provide a comprehensive survey on applications, challenges, and opportunities of Generative AI (GAI) in UAV swarms. Tian et al. [19] review the integration of LLMs with UAVs, and investigate salient tasks and application scenarios where UAVs and LLMs converge. Javaid et al. [20] summarize the state-of-the-art LLM-based UAV architectures and review opportunities for embedding LLMs within UAV frameworks, particularly in the context of enhancing data analysis and decision-making processes. Dharmalingam et al. [21] develop a framework that integrates multiple LLMs to improve the safety and operational efficiency of UAV missions. In this framework, multiple specialized LLMs are deployed for various tasks, such as inferencing, anomaly detection, and prediction, in onboard systems, edge servers, and cloud servers.
The existing literature extensively investigates the application of LLMs in wireless communications and UAV networks. However, there remains a clear gap in the use of LLM-guided scheduling and resource allocation specifically for UAVs. Furthermore, most prior studies overlook the security vulnerabilities of LLMs when deployed in autonomous systems. In contrast, this work introduces the ICLDC scheme, a novel in-context learning-based framework for UAV-assisted data collection in emergency scenarios such as SAR missions. The ICLDC optimizes the data collection schedule based on natural language task descriptions generated from logged sensory data. It continuously adapts by incorporating feedback into the task description, using this evolving context to refine future decisions. This adaptive scheduling strategy, leveraging LLMs and in-context learning, represents a significant departure from prior work. In addition, we illustrate the impact on ICLDC of jailbreaking attacks that manipulate task prompts to degrade network performance, thereby exposing and addressing a critical vulnerability of LLMs in mission-critical environments.
IV System Model
This section presents the system model of the considered UASNETs. The considered network contains $N$ ground sensors and a UAV. The UAV flies along pre-determined routes which consist of several waypoints to cover all the ground sensors in the field. The location of the UAV on its trajectory at time $t$ is denoted by $\zeta(t)$. The UAV is responsible for collecting sensory data from the ground sensors.
The coordinates $(x,y,h)$ and $(x_{j},y_{j},0)$ represent the position of the UAV and ground sensor $j$, respectively. The UAV flies to the ground sensors with constant velocity, collects sensory data, makes decisions using the proposed ICLDC.
We consider that the UAV moves in low attitude for data collection, where the probability of LoS communication between the UAV and ground sensor $j$ is given by Eq. 1 where $a$ and $b$ are constants, and $\varphi_{j}$ denotes the elevation angle between the UAV and ground sensor $j$[27].
$\Pr{{}_{LoS}}(\varphi_{j})=\frac{1}{1+a\exp(-b[\varphi_{j}-a])}$ (1)
$\varphi_{j}=\tan^{-1}\left(\frac{h}{d_{j}}\right),$ (2)
The path loss of the channel between the UAV and ground sensor $j$ is given by Eq. 3 where $r$ is the radius of the radio coverage of the UAV, $\lambda$ is the carrier frequency, and $v_{c}$ is the speed of light. $\eta_{LoS}$ and $\eta_{NLoS}$ are the excess path losses of LoS or non-LoS[28].
$\gamma_{j}=\Pr{{}_{LoS}}(\varphi_{j})(\eta_{LoS}-\eta_{NLoS})+20\log\left(r\sec(\varphi_{j})\right)+\\ 20\log(\lambda)+20\log\left(\frac{4\pi}{v_{c}}\right)+\eta_{NLoS}$ (3)
IV-A Problem Formulation
The goal is to optimize the UAV data collection schedule to minimize the total packet loss across all ground sensors. Packet loss occurs due to communication errors caused by poor channel conditions and buffer overflows caused by high queue lengths at the ground sensors. To minimize packet loss, the UAV must dynamically schedule its communication with the ground sensors based on logged environmental data (e.g., queue lengths, and channel conditions).
Let $t_{j}$ denote the time instant at which ground sensor $j$ transmits. Let $q_{j}(t)$ represent the queue length of ground sensor $j$ at time $t$, and let $D$ be the queue depth. Additionally, let $\gamma_{j}$ be the channel gain of ground sensor $j$, and $\gamma_{th}$ be the channel gain threshold. The optimization of the UAV communication schedule aims to minimize the packet loss of all the ground sensors, as given by Eq. 4 with constraints expressed in Eqs. 4a and 4b.
Figure 2: Data communication protocol for the proposed ICLDC. ICLDC is conducted at the UAV where in each communication frame queries the LLM to determine the sensor selection.
$\min_{t_{j},j=1..N}$ $\quad\sum_{j=1}^{N}\Big{(}f_{j}(t_{j}))+g_{j}(t_{j})\Big{)}$ (4)
$f_{j}(t)$ =$=\begin{cases}1,&\text{if }(t=t_{j})\text{ and }(\gamma_{j}\leq\gamma_{th}),\\ 0,&\text{otherwise}.\end{cases}$ (4a)
$g_{j}(t)$ =$=\begin{cases}1,&\text{if }(q_{j}(t)>D)\text{ and }(t\neq t_{j}),\\ 0,&\text{otherwise}.\end{cases}$ (4b)
IV-B Communication Protocol
The proposed communication protocol for data collection in UASNETs using the proposed ICLDC is depicted in Fig. 2. The UAV queries the edge LLM to select the most relevant sensor for data collection $s_{j}$ based on queue length $q_{j}$, and channel conditions $\gamma_{j}$. Once the sensor is selected and the UAV is positioned at an adequate distance, the UAV sends a beacon message containing the sensor ID. The selected sensor sends its data packets, including sensor data and status information (queue length and channel conditions). The UAV checks the received data and sends a confirmation to the sensor. Then the UAV can move to a following sensor and repeat the operation. After a certain time, a new time step (episode) begins, with the LLM being queried again for an updated sensor schedule, and the data collection continues.
@@ IV-B Communication Protocol
Table II: Summary of Jailbreaking in LLMs: Attack strategies used in this work

### v2

@@ Abstract
Unmanned Aerial Vehicles (UAVs) are increasingly being utilized in various private and commercial applications, e.g., traffic control, parcel delivery, and Search and Rescue (SAR) missions. Machine Learning (ML) methods used in UAV-Assisted Sensor Networks (UASNETs) and, especially, in Deep Reinforcement Learning (DRL) face challenges such as complex and lengthy model training, gaps between simulation and reality, and low sampling efficiency, which conflict with the urgency of emergencies, such as SAR missions. In this paper, an In-Context Learning (ICL)-Data Collection Scheduling (ICLDC) system is proposed as an alternative to DRL in emergencies. The UAV collects sensory data and transmits it to a Large Language Model (LLM), which creates a task description in natural language. From this description, the UAV receives a data collection schedule that must be executed. A verifier ensures safe UAV operations by evaluating the schedules generated by the LLM and overriding unsafe schedules based on predefined rules. The system continuously adapts by incorporating feedback into the task descriptions and using this for future decisions. This method is tested against jailbreaking attacks, where the task description is manipulated to undermine network performance, highlighting the vulnerability of LLMs to such attacks. The proposed ICLDC significantly reduces cumulative packet loss compared to both the DQN and Maximum Channel Gain baselines. ICLDC presents a promising direction for intelligent scheduling and control in UASNETs.
@@
Unmanned Aerial Vehicles, Large Language Models, In-Context Learning, Network Edge, Jailbreaking Attacks
@@ I Introduction
Unmanned Aerial Vehicles (UAVs) enjoy controlled mobility and rapid deployment. These features allow them to be used in many civilian and commercial applications, e.g., traffic control[1], parcel delivery[2], and crop monitoring[3]. UAVs are also employed to collect sensory data in harsh environments, such as natural disaster monitoring, and border surveillance [4]. In particular, Search and Rescue (SAR) missions are critical endeavors in disaster management and public safety, often involving high-risk situations where human lives are at stake. In recent years, the integration of UAV technology into SAR missions has revolutionized the field, offering a powerful tool to enhance the efficiency, effectiveness, and safety of rescue missions [5]. A common trend in SAR missions is the use of Deep Reinforcement Learning (DRL), as demonstrated by Peake et al.[6] and Kurunathan et al.[7], to optimize UAV-Assisted Sensor Networks (UASNETs) in disaster areas. However, DRL optimizers often suffer from low sample efficiency, simulation-to-reality gaps, and the need for complex model training and fine-tuning, which are not compatible with the urgent and emergency nature of SAR missions[8],[9].
@@
On the other hand, Large Language Models (LLMs) are transformative technologies with successful applications in education, finance, healthcare, biology, and more [10]. Their potential extends to network management and optimization, particularly through In-Context Learning (ICL), a feature that allows LLMs to learn from language-based descriptions and demonstrations. Compared to existing methods like convex optimization or conventional Machine Learning (ML), LLM-assisted ICL offers key advantages. It relies on the LLM’s inference process, eliminating the need for complex model training and fine-tuning, which are often bottlenecks for many existing ML techniques. In addition, tasks can be formulated using natural language, making it accessible to operators without professional technical expertise.
@@
These benefits make LLM-assisted ICL a promising approach for simplifying and enhancing network management and optimization tasks[11]. For instance, UAVs can use LLM-assisted ICL to optimize data collection scheduling and dynamically adapt to changing environments. However, the use of LLMs also introduces vulnerabilities to jailbreaking attacks, which refer to attempts to bypass or undermine the safety filters and restrictions integrated into LLMs. The main objective is to circumvent the model’s built-in limitations, enabling it to perform actions or produce outputs that are typically restricted or prohibited.
Figure 1: Overall design of the proposed ICLDC. The UASNETs interact with an edge-hosted LLM through structured prompts. The LLM receives logged environmental data (e.g., queue length, channel conditions, and battery level) and generates optimized data collection schedules. A feedback loop records system performance and guides continuous adaptation to minimize packet loss.
@@
Given the challenges and opportunities of LLMs, this paper proposes ICL-based Data Collection Scheduling (ICLDC) scheme. Specifically, as depicted in Fig. 1, we conduct ICL to optimize the data collection schedule in UASNETs. The UAV interacts with the environment to collect logged sensory data and transmit it to the LLM for analysis and decision-making. The LLM is initialized with a robust system prompt that defines its role, constraints, and ethical guidelines, ensuring it operates safely and securely. In particular, the LLM begins by crafting a detailed task description in natural language, which includes the task goal, input data, rules, examples, output, and feedback. Then, the LLM leverages this information to generate a data collection schedule, which the UAV then executes. In addition, a verifier is designed to validate the data collection schedule generated by the LLM. The verifier employs a scoring system based on critical safety rules, such as queue length, channel condition, and battery level, to evaluate the LLM’s selection. Suppose the LLM chooses a sensor that violates safety principles (i.e., receives a low safety score). In that case, the verifier overrides the decision and selects the sensor with the highest score, ensuring safe and logical actions. Following this, the data collection schedule made by the LLM, along with the environment state (e.g., queue length, battery level, and channel condition of ground sensors) and the resulting system performance, such as packet loss, is recorded as feedback. This feedback is appended to the task description and included in subsequent queries to the LLM. By incorporating feedback, the LLM can avoid repeating past mistakes (e.g., selecting sensors with poor channel conditions) and improve its future decision-making. Under attack, the task description is manipulated to negatively impact the network cost. ICLDC detects attacks by measuring prompt perplexity to flag suspicious LLM requests and dynamically computes attack probability to assess threats. Epsilon is adjusted based on attack probability, increasing randomness (exploration) when risks escalate to prioritize security. In high-risk cases, it executes randomized safe actions, ensuring robust defense against adversarial inputs.
@@
  • ICLDC, an LLM-assisted ICL to optimize the data collection schedule in UASNETs. In contrast to conventional methods such as DRL, which require extensive model training and fine-tuning, the proposed ICLDC enables the UAV to schedule data collection based on natural language descriptions and demonstrations. This approach is instrumental in dynamic environments such as disaster monitoring or SAR missions.
@@
The rest of this paper is organized as follows: Section II discusses background concepts on UAV data collection, ICL, jailbreaking attacks, and network edge. Section III presents the literature review. Section IV presents the system model comprising problem formulation and communication protocol. Section V presents the attack model with a detailed procedure. Section VI presents the proposed approach. Section VII is devoted to numerical results and discussions, and finally, Section VIII concludes the paper.
@@ II Background Concepts
This section introduces some background concepts related to UAV data collection, ICL, edge and jailbreaking, including attack types and template completion attacks.
@@
Table I: Summary of Jailbreaking in LLMs: Attack strategies used in this work
@@
II-A UAV Data Collection
UASNETs leverage the unique capabilities of UAVs to operate in SAR missions and remote environments, making them ideal for aerial data collection. UAVs can serve as Aerial Base Stations (ABS) or relays, extending the coverage and connectivity of sensor networks. In UASNETs, the UAV is employed to hover over the target area, moving sufficiently close to ground sensors, exploiting short-distance LoS communication links for data collection. However, selecting a ground sensor for data collection may lead to buffer overflows at other sensors if those sensors’ buffers are already full while new data keeps arriving. Moreover, the transmissions of ground sensors that are far away from the UAV and experience poor channel conditions are prone to errors at the UAV. Adequately scheduling data collection is critical to prevent data queue overflow and communication failure[12],[13].
II-B In-Context Learning
II-C Jailbreaking Attacks
A prominent black-box method, template completion attacks, disguises harmful queries within structured templates, compelling LLMs to complete dangerous responses unknowingly. Attackers achieve this through scenario nesting (framing queries within fictional or role-playing contexts), context-based manipulation (using misleading continuity to elicit unauthorized outputs), and code injection (embedding malicious instructions within programming tasks). These attacks are particularly dangerous as they evade keyword-based filters, exploit the LLM’s pattern recognition capabilities, and work effectively even against well-aligned models[15]. Table I provides a summary of jailbreaking attacks in LLMs.
II-D Edge
Deploying LLMs at the edge for UAV applications requires optimizing large models through techniques like quantization, pruning, and parameter-efficient fine-tuning methods, which significantly reduce memory and computational demands to fit limited onboard hardware. This approach enables real-time inference for UAVs that demand near-instantaneous decision-making during autonomous operations.
Because the data is processed directly on the device, Edge AI enables fast responses that are crucial for applications such as autonomous vehicles. By moving computation to the edge, latency is significantly reduced, making this approach ideal for low-latency applications such as data collection using UAVs. In this context, a framework for ubiquitous distributed wireless edge computing to support real-time LLM inference is proposed in [16]. In particular, they present a novel LLM edge inference framework that incorporates batching and model quantization to achieve high-throughput inference on resource-constrained edge devices. Additionally, a comprehensive system is proposed that intelligently adapts LLMs for efficient, latency-aware inference on commercial off-the-shelf edge hardware [17]. The system works in two key phases: (1) offline device-specific model adaptation and (2) online latency-aware system optimization.
Table II: GPT-4o-mini Architecture Specifications
Architecture Aspect GPT-4o-mini Specification
Base Design Transformer (decoder-only)
Size Scaled-down from GPT-4 Turbo (exact parameters undisclosed)
Context 128K tokens ($\approx$300-400 pages) with efficient attention mechanisms
Efficiency Likely uses Mixture-of-Experts (MoE) with expert pruning
Training Pretrained on diverse data up to 2023, RLHF-tuned
Inference FP16/FP8 precision $\bullet$ KV caching $\bullet$ Flash Attention optimization
Tokenization Byte-pair encoding (BPE)
Throughput Optimized for faster response times vs. larger GPT-4 variants
Special Features Supports long-document processing $\bullet$ JSON mode $\bullet$ better cost-performance tradeoff
5G networks, with mmWave communication, massive MIMO, and beamforming, enable Ultra-Reliable Low-Latency Communication (URLLC), ideal for mission-critical applications like UAV-Assisted SAR. URLLC achieves end-to-end latencies as low as 1 millisecond, supporting real-time decision-making in time-sensitive scenarios. These high-throughput, low-latency links between UAVs and edge infrastructure allow rapid transmission of sensor data to edge devices. This facilitates on-the-fly LLM processing, enabling responsive and context-aware decisions in dynamic environments.
While lightweight models like OPT-125M are deployable on embedded UAV hardware [18], they lack the reasoning capacity for complex natural language tasks. More capable open-source models (e.g., LLaMA-2, Phi-2) are technically challenging to integrate in UAV environments due to runtime and infrastructure constraints. Proprietary models such as GPT-4o-mini and GPT-3.5-Turbo, despite requiring external APIs, offer practical advantages in terms of ease of deployment, reproducibility, and support for advanced language understanding necessary for the study’s use case.
The scalability of LLMs at the edge is fundamentally constrained by three interrelated factors: memory capacity, computational throughput, and energy efficiency, all of which are significantly more limited on edge nodes than in cloud environments. The theoretical upper limit of the LLM scale on edge nodes can be estimated based on available system memory and model quantization. Given that LLM weights must reside in memory, the total parameter capacity P is approximately:
$P=\frac{M\times 8}{b}$ (1)
where M is the available memory in bytes, and b is the bit precision per parameter (e.g., 16 for FP16, 8 or 4 for quantized models). For example, an edge device with 8–16 GB of usable RAM and FP16 precision could support 4–8 billion parameters. With 4-bit quantization, this upper bound increases to approximately 16–32 billion parameters. However, accounting for runtime overhead and activations, the practical range is significantly lower, often capped at 1 to 10 billion parameters, depending on the level of compression and workload. This aligns with the current industry trends: models such as Gemini Nano [19] (1.8B–3.25B parameters) and Qualcomm’s on-device LLMs [20] ($\leq$ 10B) represent the high end of what is practically feasible on advanced edge hardware. Thus, the theoretical upper limit of LLM scale that a single edge node can support is on the order of $^{9}1010$$^{10}$ parameters, with practical deployments today typically using models in the 1–3 billion parameter range. Beyond this, cloud offloading or distributed execution becomes necessary [21], [22].
Table II summarizes the internal structure of the LLM deployed at the edge. The table summarizes the core architectural aspects of GPT-4o-mini, a streamlined variant of GPT-4 Turbo designed for efficiency and long-context handling. It is built on a decoder-only Transformer architecture. The model supports a context window of up to 128K tokens—equivalent to approximately 300–400 pages—using efficient attention mechanisms. It incorporates a Mixture-of-Experts approach with expert pruning, while its inference is further optimized through FP16/FP8 precision, KV caching, and Flash Attention. It employs byte-pair encoding for tokenization and is tuned for higher throughput compared to larger GPT-4 models.
III Related Work
When searching the current literature for the use of LLM-assisted ICL in the design, deployment, and operation of wireless systems and UASNETs, we realized that this topic is rarely investigated, particularly in the context of UASNETs and security. This section briefly mentions the related work we found and its focus.
Zhou et al. [11] propose an LLM-assisted ICL algorithm for base station transmission power control that can handle both discrete and continuous state problems; their approach outperforms state-of-the-art DRL algorithms. In another work, Zhou et al. [14] provide a detailed investigation of different prompting techniques such as ICL, chain-of-thought, and self-refinement, and propose novel prompting schemes for network optimization, and the case studies confirm their satisfactory performance. Dong et al. [23] present a comprehensive survey on ICL where advanced techniques, comprising training and prompt designing strategies, are discussed. Zheng et al. [24] investigate the potential of deploying LLMs at the 6G edge, identify the critical challenges for LLM deployment at the edge, and envision the 6G Mobile Edge Computing (MEC) architecture for LLMs. Liu et al. [25] provide a comprehensive survey on applications, challenges, and opportunities of Generative AI (GAI) in UAV swarms. Tian et al. [26] review the integration of LLMs with UAVs and investigate salient tasks and application scenarios where UAVs and LLMs converge. Javaid et al. [27] summarize the state-of-the-art LLM-based UAV architectures and review opportunities for embedding LLMs within UAV frameworks, particularly in the context of enhancing data analysis and decision-making processes. Dharmalingam et al. [28] develop a framework that integrates multiple LLMs to improve the safety and operational efficiency of UAV missions. In this framework, multiple specialized LLMs are deployed for various tasks, such as inferencing, anomaly detection, and prediction, in onboard systems, edge servers, and cloud servers.
Figure 2: The proposed ICLDC uses a data communication protocol where each communication frame queries the LLM to select the sensor.
The existing literature extensively investigates the application of LLMs in wireless communications and UAV networks. However, there remains a clear gap in the use of LLM-guided scheduling and resource allocation specifically for UASNETs. Moreover, most prior studies overlook the security vulnerabilities of LLMs when deployed in autonomous systems. In contrast to [14] and [11], which focus on simplified power control with single-variable environments (e.g., user distance or number), our work tackles a more challenging UAV scheduling problem, jointly optimizing queue lengths, battery levels, and channel conditions, providing a rigorous evaluation of in-context learning. Moreover, we rigorously evaluate ICL under adversarial attacks, a critical gap in [14] and [11], ensuring practical viability for real-world deployments. This work introduces the ICLDC scheme, a novel ICL-based framework for UASNETs in emergency scenarios such as SAR missions. The ICLDC optimizes the data collection schedule based on natural language task descriptions generated from logged sensory data. It continuously adapts by incorporating feedback into the task description, using this evolving context to refine future decisions. This adaptive scheduling strategy, leveraging LLMs and ICL, represents a significant departure from prior work. In addition, we illustrate the impact on ICLDC of jailbreaking attacks that manipulate task prompts to degrade network performance, thereby exposing and addressing a critical vulnerability of LLMs in mission-critical environments.
IV System Model
This section presents the system model of the considered UASNETs. The considered network contains $N$ ground sensors and a UAV. The UAV flies along pre-determined routes which consist of several waypoints to cover all the ground sensors in the field. The proposed ICLDC assumes that the UAV trajectories are predefined to be consistent with previous studies aimed at improving network capacity [34], increasing coverage [35], or reducing propulsion energy consumption [36]. Pre-defining the flight path also avoids the additional energy costs and complexity associated with real-time flight path planning. Given these considerations, we focus on the data collection schedule to minimize data loss due to buffer overflows and channel fading. Importantly, the proposed ICLDC framework is generic and can be applied to any UAV trajectory. The location of the UAV on its trajectory at time $t$ is denoted by $\zeta(t)$. The UAV is responsible for collecting sensory data from the ground sensors.
The coordinates $(x,y,h)$ and $(x_{j},y_{j},0)$ represent the position of the UAV and ground sensor $j$, respectively. The UAV flies to the ground sensors with constant velocity, collects sensory data, and makes decisions using the proposed ICLDC.
We consider that the UAV moves in low altitude for data collection, where the probability of LoS communication between the UAV and ground sensor $j$ is given by Eq. 2 where $a$ and $b$ are positive constants, and $\varphi_{j}$ denotes the elevation angle between the UAV and ground sensor jj[37].
$P_{LoS}(\varphi_{j})=\frac{1}{1+a\exp(-b(\varphi_{j}-a))}$ (2)
$\varphi_{j}=\tan^{-1}\left(\frac{h}{d_{j}}\right),$ (3)
The path loss of the channel between the UAV and ground sensor $j$ is given by Eq. IV where $r$ is the radius of the radio coverage of the UAV, $\lambda$ is the carrier frequency, and $v_{c}$ is the speed of light. $\eta_{LoS}$ and $\eta_{NLoS}$ are the excess path losses of LoS or non-LoS[38].
$\gamma_{j}=$ $P_{LoS}(\varphi_{j})(\eta_{LoS}-\eta_{NLoS})+20\log\left(r\sec(\varphi_{j})\right)+$
$20\log(\lambda)+20\log\left(\frac{4\pi}{v_{c}}\right)+\eta_{NLoS}$ (4)
Table III: Summary of the Attack Model (LLM: Large Language Model; UAV: Unmanned Aerial Vehicle).
Impact Increased packet loss [%].
IV-A Problem Formulation
The goal is to optimize the UAV data collection schedule to minimize the total packet loss across all ground sensors. Packet loss occurs due to communication errors caused by poor channel conditions and buffer overflows caused by excessive queuing at the ground sensors. To minimize packet loss, the UAV must dynamically schedule its communication with the ground sensors based on logged environmental data (e.g., battery level, queue lengths, and channel conditions).
Let $t_{j}$ denote the time instant at which ground sensor $j$ transmits. Let $q_{j}(t)$ represent the queue length of ground sensor $j$ at time $t$, and let $D$ be the buffer depth. Additionally, let $\gamma_{j}$ be the channel gain of ground sensor $j$, and $\gamma_{th}$ be the channel gain threshold.
The UAV communication schedule consists of the set of time instants in which the UAV visits all sensors once [^1] For convenience, in the remainder of the paper we assume that the communications schedule corresponds to a sequence of sensors from which the visit times can be computed.. The schedule optimization aims to minimize the packet loss of all the ground sensors in a full data gathering cycle, as given by Eq. 5 with constraints expressed in Eqs. 5b and 5c.
$\min_{\{t_{j},j=1..N\}}$ $\sum_{t_{j}=t_{1}}^{t_{N}}\sum_{i=1}^{N}\left(f_{i}(t_{j})+g_{i}(t_{j})\right)$ (5a)
$f_{i}(t_{j})$ $=\begin{cases}1,&\text{if }(t_{i}=t_{j})\text{ and }(\gamma_{j}\leq\gamma_{th}),\\ 0,&\text{otherwise},\end{cases}$ (5b)
$g_{i}(t_{j})$ $=\begin{cases}1,&\text{if }(q_{i}(t_{j})>D)\text{ and }(t_{i}\neq t_{j}),\\ 0,&\text{otherwise}.\end{cases}$ (5c)
IV-B Communication Protocol
The proposed communication protocol for data collection in UASNETs using the proposed ICLDC is depicted in Fig. 2. The UAV queries the edge LLM to select the most relevant sensor $j$ for data collection based on battery level $b_{j}$, queue length $q_{j}$, and channel conditions $\gamma_{j}$. Once the sensor is selected and the UAV is positioned at an adequate distance, the UAV sends a beacon message containing the sensor ID. The selected sensor sends its data packets, including sensor data and status information (battery level, queue length, and channel conditions). The UAV checks the received data and sends a confirmation to the sensor. Then the UAV can move to the following sensor and repeat the operation. After a certain time, a new time step begins, with the LLM being queried again for an updated sensor schedule, and the data collection continues.